OpenStudy (anonymous):

four more than 6 times a number is the same as 9 times the number increased by 10.

OpenStudy (gk99):

4 + 6n = 9n + 10 -3n = 6 n = -2

OpenStudy (anonymous):

4+6x=9x+10